A particle moves in x-y plane. At time t=0, particle is at (1m,2m) and has vel.(4i+6j)m/s. At t=4sec, particle  reaches (6m,4m) and has velocity (2i+10j)m/s. In the given time interval, find:


a)Avge.Vel=?


b)Avge. Acceleration=?


c)From,given data, find avge.Speed=?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chilukuri

Total displace ment = ( (6-1)2 + ( 4-2)2 )1/2

                                            = √29  m

total time     = 4 sec

change in velocity = |(2i+10j) -(4i +6j) |  =√20

 

so avg velocity = total displacement /total time

                     =√29 /4  m/sec

avg acceleration      = change in velocity /time interval

                            = √20 /4  =√5 /2  m/sec2

since path of the particle is not know only initial and final position is known so avg speed can not be calculated with this given data


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin